OpenStudy (anonymous):

The formula for the volume of a sphere is . Find the radius, to the nearest hundredth, of a sphere with a volume of 15 in.3. A. 3.58 in. B. 258.01 in. C. 1.53 in. D. 1.85 in.

OpenStudy (anonymous):

@a

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@adl_thoo

OpenStudy (anonymous):

@BulletWithButterflyWings

OpenStudy (anonymous):

@.Sam.

OpenStudy (kropot72):

the formula for the volume of a sphere is \[V=\frac{4}{3} \pi r ^{3}\] \[15=\frac{4}{3} \pi r ^{3}\] If you divide both sides of the equation by \[\frac{4}{3} \pi\] you will have an equation for r cubed. Can you do that?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

how do u do it I don't know

OpenStudy (kropot72):

\[r ^{3}=15\times \frac{3}{4 \pi}=?\]

OpenStudy (anonymous):